the transformer for a particular device can produce 6.50 V and a maximum current of 2.10 A. At this maximum value, find the current drawn when the transformer is connected to a 120 V input
Answer #1

Considering the Volt-Ampere balance between the input and output of the transformer,

V1 I1 = V2 I2

(6.50)*(2.10) = (120) * I2

I2 = ( 6.50 * 2.10 ) / (120) = 0.11375 A = 11.375 mA (Ans)